OpenStudy (anonymous):

IF 5.4 J of work are needed to stretch a spring from 17 cm to 23 cm and 9 J are needed to stretch it from 23 cm to 29cm, what is the natural length of the spring.

OpenStudy (anonymous):

I know Hooke's Law and I got \[\int\limits_{0}^{.06} kxdx=5.4\] which turns out to be k=30 but I'm not sure where to go from here.

OpenStudy (dumbcow):

i have a theory ...do you know the answer by chance?

OpenStudy (anonymous):

no its for homework

OpenStudy (dumbcow):

ok in Hooks Law F = kx , the "x" is distance from natural length so \[5.4 = k(x+6)\] \[9 = k(x+12)\] solving the system \[k = 0.6, x = 3\] natural length of spring is 14 cm

OpenStudy (dumbcow):

get 2nd opinion though, not expert on physics stuff

OpenStudy (anonymous):

im gonna go with it. Seems right to me. p.s. the reason I didn't ask you this question was to give you a break from me haha

OpenStudy (dumbcow):

haha

OpenStudy (anonymous):

buttt since you are here find the volume generated by rotating the given region about the specified line y=6x y=0 x=1

OpenStudy (anonymous):

 I know this is my graph

OpenStudy (anonymous):

is my integral \[\int\limits_{0}^{6}\pi 6x\]

OpenStudy (dumbcow):

what is it rotating around? x=1 ?

OpenStudy (anonymous):

This is the graph actually and the question says find the volume generated by rotating the given region about the specified line. R1 about AB so yea i think its 1

OpenStudy (dumbcow):

oh ok , well since region is just a nice triangle no calculus needed, essentially you are finding volume of a cone

OpenStudy (anonymous):

a cone is 1/2 pi r^2 right?

OpenStudy (anonymous):

my r being 6x

OpenStudy (dumbcow):

close \[V = \frac{1}{3}\pi r^{2}h\]

OpenStudy (anonymous):

oh so my h is actually 6x and r is x right?

OpenStudy (dumbcow):

"r" is radius of circle, look at base yes now you got it

OpenStudy (dumbcow):

for cone r = 1 h = 6 the set up using shell method is: \[2\pi \int\limits_0^1 x(6x) dx\]

OpenStudy (anonymous):

i got 4pi

OpenStudy (dumbcow):

oops my radius is off....r = 1-x using volume of cone, you get 2pi

OpenStudy (anonymous):

ohhh wait why 1-x?

OpenStudy (dumbcow):

radius is distance from axis of rotation (x=1) so when x=1, the radius should be 0 when x=0, radius should be 1

OpenStudy (anonymous):

oh ok i see

OpenStudy (anonymous):

so yea i got the 2pi

OpenStudy (anonymous):

ok one more and then I'm off to bed and I will leave you alone. Till tomorrow ;) Find the volume V generated by rotating the given region about the specified line. R3 about OC http://www.webassign.net/waplots/c/7/5b24e5fe28d5d47788e1c5399cf12f.gif

OpenStudy (anonymous):

I wanna use the cylindrical shell method so i know i will have \[2\pi \int\limits_{?}^{?}rhdx\]

OpenStudy (dumbcow):

yes good do you know what the limits will be?

OpenStudy (anonymous):

I wanna say from 0 to 3

OpenStudy (dumbcow):

integrating wrt "x" not "y"

OpenStudy (anonymous):

but its around the y=axis

OpenStudy (dumbcow):

true but we are using shell method if using disc method then in terms of "y"

OpenStudy (anonymous):

so shells and disc method are not the same limits?

OpenStudy (dumbcow):

no disc method you are summing up circle cross-sections perpendicular to axis shell method you sum up surface areas parallel to axis that site i sent you has some visual diagrams illustrating both methods

OpenStudy (anonymous):

ok i was just asking bc we had another homework question that was just like this and they said do both limits and i used the same limits

OpenStudy (anonymous):

do both methods*

OpenStudy (dumbcow):

anyway, look at radius, it goes out horizontally from vertical axis that is clue everything in terms of "x"

OpenStudy (anonymous):

ok so from 0 to 1

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

you are looking at the yellow part of the graph correct?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

ok just checking lol

OpenStudy (anonymous):

\[2\pi \int\limits_{0}^{1} x(3\sqrt[4]{x} -3x)\]

OpenStudy (dumbcow):